# Products of idempotents in a quaternion ring

### v1

@@ 1. Introduction
In this paper, we study products of idempotents in quaternion rings over finite commutative local principal rings. Our first main result shows that, in this setting, allowing arbitrarily many idempotent factors does not enlarge the class of representable elements: every element of a quaternion ring that is a product of idempotents is already a product of two idempotents (Theorem 3.4). Using this characterization, we then determine the exact number of elements of a quaternion ring over a finite commutative local principal ring that admit such a factorization (Theorem 3.8).
@@ 2. Definitions and preliminaries
Let R be a finite local commutative ring. If $A\in M_{2}(R)$ is a nontrivial idempotent, then $A\in\mathcal{O}_{M(1,0)}$.
@@ 3. Products of idempotents
In this section we turn to the product of idempotents in a quaternion ring over a finite commutative local ring. We begin by determining the number of idempotent elements in such rings.
@@
Suppose firstly that $p>2$. By [cher, Theorem 3.10], we know that $H(R)\simeq M_{2}(R)$. Lemmas 2.2 and 2.3 from [dolzan] now give us $|I(H(R))|=2+q^{3n-2}(q^{2}-1)$. On the other hand, if $p=2$, 22 is a zero-divisor in $R$, so $H(R)$ is a finite local ring by [cherdol, Lemma 3.1] and as such contains no nontrivial idempotents. ∎
@@
We may therefore assume that $2\in R$ is invertible. Then we have $H(R)\simeq M_{2}(R)$ by [cher, Theorem 3.10], so we have to prove that any $A\in M_{2}(R)$ that is a product of $r$ idempotent matrices for some $r\geq 3$, can also be written as a product of two idempotents. Therefore, assume that $A=E_{1}E_{2}\ldots E_{r}$ for idempotent matrices $E_{1},E_{2},\ldots,E_{r}$. We can also assume without any loss of generality that $E_{1}$ is a nontrivial idempotent. By Lemma 2.1, there exists an invertible matrix $P\in M_{2}(R)$ such that $PE_{1}P^{-1}=M(1,0)$. Thus $PAP^{-1}=M(1,0)F_{2}\ldots F_{r}$ for some (idempotent) matrices $F_{2},\ldots,F_{r}$. This implies that $PAP^{-1}=M(a,b)$ for some $a,b\in R$. By Lemma 3.3, this implies that $A$ is a product of two idempotents. ∎
@@
[…]
@@
We know that $\left|\mathcal{O}_{M(a,b)}\right|=\frac{|GL_{2}(R)|}{|\mathrm{Stab}(M(a,b))|}$, so we have to calculate the sizes of the respective stabilizers.
Suppose firstly that $a=0$ and $b\neq 0$. Observe that $P=\left(\begin{array}[]{cc}\alpha&\beta\\ \gamma&\delta\end{array}\right)\in\mathrm{Stab}(M(0,b))$ if and only if $PM(0,b)=M(0,b)P$ if and only if $b\gamma=b(\alpha-\delta)=0$. Since $b\neq 0$, we have $\gamma\in J(R)^{n-k}\subseteq J(R)$, so $\alpha,\delta\in U(R)$ by the fact that $P$ is an invertible matrix. By Lemma 2.2, $|J(R)^{n-k}|=q^{k}$ and $|U(R)|=q^{n}-q^{n-1}=q^{n-1}(q-1)$, so we have $q^{n+k-1}(q-1)$ possibilities for choosing $\gamma$ and $\alpha$. Since $\beta$ is arbitrary ($q^{n}$ possibilities) and $\delta\in\alpha+J(R)^{n-k}$ ($q^{k}$ possibilities), we have $|\mathrm{Stab}(M(0,b))|=q^{2n+2k-1}(q-1)$. We know that $|GL_{2}(R)|=q^{4n-3}(q-1)(q^{2}-1)$ (see for example [dolzan, Lemma 2.3]), therefore we have $\left|\mathcal{O}_{M(0,b)}\right|=q^{2n-2k-2}(q^{2}-1)$.
We can assume henceforth that $a\neq 0$. Let us continue with calculating $|\mathrm{Stab}(M(a,0))|$. Observe that $P=\left(\begin{array}[]{cc}\alpha&\beta\\ \gamma&\delta\end{array}\right)\in\mathrm{Stab}(M(a,0))$ if and only if $\gamma a=\beta a=0$, so $\beta,\gamma\in J(R)^{n-l}$. Therefore, we have $q^{2l}$ possibilities for choosing $\beta$ and $\gamma$. Since $a\neq 0$, we have $\beta,\gamma\in J(R)$, so $\alpha,\delta\in U(R)$. This gives us $|\mathrm{Stab}(M(a,0))|=q^{2(l+n-1)}(q-1)^{2}$ and therefore $\left|\mathcal{O}_{M(a,0)}\right|=q^{2n-2l-1}(q+1)$.
Let us finally calculate $|\mathrm{Stab}(M(a,b))|$ in the case $k<l<n$. Observe that $P=\left(\begin{array}[]{cc}\alpha&\beta\\ \gamma&\delta\end{array}\right)\in\mathrm{Stab}(M(a,b))$ if and only if $PM(a,b)=M(a,b)P$ if and only if $\gamma a=\gamma b=0$ and $a\beta=(\alpha-\delta)b$. Since $k<l$, the conditions $\gamma a=\gamma b=0$ are equivalent to the fact that $\gamma\in J(R)^{n-k}$. Since $|J(R)^{n-k}|=q^{k}$, we have $q^{k}$ possibilities for choosing $\gamma$. By Lemma 2.2, there also exists $x\in J(R)$ such that $J(R)^{t}=(x^{t})$ for every $t=0,1,\ldots,n$. Therefore $a=ux^{l}$ and $b=vx^{k}$ for some $u,v\in U(R)$. Since $\gamma\in J(R)$, the fact that $P$ is an invertible matrix implies that $\alpha,\delta\in U(R)$. Choose any $\beta\in R$ and any $\alpha\in U(R)$, giving us $q^{n}$ and $q^{n-1}(q-1)$ possibilities, respectively. The condition $a\beta=(\alpha-\delta)b$ can be rewritten as $v^{-1}u\beta x^{l}=(\alpha-\delta)x^{k}$. Denote $v^{-1}u\beta=\beta_{0}+\beta_{1}x+\ldots+\beta_{n-1}x^{n-1}$ and $\alpha-\delta=\epsilon_{0}+\epsilon_{1}x+\ldots+\epsilon_{n-1}x^{n-1}$. This implies $\epsilon_{0}=\epsilon_{1}=\ldots=\epsilon_{l-k-1}=0$ and $\epsilon_{l-k}=\beta_{0},\epsilon_{l-k+1}=\beta_{1},\ldots,\epsilon_{n-k-1}=\beta_{n-l}$. (Note here that $k<l$ implies that we always have $\epsilon_{0}=0$, so $\delta$ is automatically invertible if $\alpha$ is invertible.) Therefore, the $k$ elements $\epsilon_{n-k},\epsilon_{n-k+1},\ldots,\epsilon_{n-1}$ can be chosen arbitrarily (giving us $q$ choices for each one, since $J(R)^{m}/J(R)^{m+1}$ is a one-dimensional vector space over $R/J(R)$ by Lemma 2.2). Therefore, we have $|\mathrm{Stab}(M(a,b))|=q^{k}q^{n}q^{n-1}(q-1)q^{k}=q^{2k+2n-1}(q-1)$ and thus $\left|\mathcal{O}_{M(a,b)}\right|=q^{2(n-k-1)}(q^{2}-1)$. ∎
@@
We can therefore assume that $2\in R$ is invertible. Then we have $H(R)\simeq M_{2}(R)$ by [cher, Theorem 3.10], so Lemma 3.3 and Theorem 3.4 tell us that $|Z|=|\bigcup_{a,b\in R}\mathcal{O}_{M(a,b)}|+1$. Again using Lemma 2.2, we know there exists $x\in J(R)$ such that $J(R)^{t}=(x^{t})$ for every $t=0,1,\ldots,n$. Suppose $a\in J(R)^{l}\setminus J(R)^{l+1}$ and $b\in J(R)^{k}\setminus J(R)^{k+1}$ for some integer $0\leq l\leq n-1$ and some integer $0\leq k\leq n$. Using Corollary 3.7, we therefore have $\mathcal{O}_{M(a,b)}=\mathcal{O}_{M(a,0)}$ if $k\geq l$ and $\mathcal{O}_{M(a,b)}=\mathcal{O}_{M(a,x^{k})}$ if $k<l$. This gives us
@@
$|Z|=2+\sum_{k=0}^{n-1}q^{2(n-k-1)}(q^{2}-1)+\\ \sum_{l=0}^{n-1}\left(q^{n-1-l}(q-1)\left(q^{2n-2l-1}(q+1)+\sum_{k=0}^{l-1}q^{2n-k-l-1}(q^{2}-1)\right)\right).$
@@
$|Z|=2+(q^{2n}-1)+\frac{q(q^{3n}-1)(q^{2}-1)}{q^{3}-1}+q^{n+1}(q^{2(n-1)}-1)-\frac{q^{2}(q^{3(n-1)}-1)(q^{2}-1)}{q^{3}-1}=\\ 1+q^{2n}+q^{n+1}(q^{2(n-1)}-1)+\frac{q(q^{2}-1)(q^{3n}-q^{3n-2}+q-1)}{q^{3}-1},$
@@
$\frac{15\alpha^{3}+13\alpha^{2}-39\alpha+37}{13}$

### v2

@@ Abstract.
Key words and phrases:
finite ring, idempotent, quaternion, matrix
@@ 1. Introduction
In this paper, we study products of idempotents in quaternion rings over finite commutative local principal rings. Our first main result shows that in this setting, the set of elements representable as a product of idempotents coincides with the set of products of two idempotents (Theorem 3.4). Using this characterization, we then determine the exact number of elements of a quaternion ring over a finite commutative local principal ring that admit such a factorization (Theorem 3.8).
@@ 2. Definitions and preliminaries
Let $R$ be a finite commutative local ring. If $A\in M_{2}(R)$ is a nontrivial idempotent, then $A\in\mathcal{O}_{M(1,0)}$.
@@
We also have the following theorem that can be deduced from [cher, Theorem 3.10], but we include the proof here for the sake of completeness.
Theorem 2.4.
Let $R$ be a finite commutative local ring of order $p^{nr}$ such that $R/J(R)$ is a field with $p^{r}$ elements, for some odd prime number $p$ and integers $n,r$. Then $H(R)\simeq M_{2}(R)$.
Proof.
By the proof of [cher, Theorem 3.10], we only need to prove that there exist $\alpha,\beta\in R$ such that $1+\alpha^{2}+\beta^{2}=0$. We prove this by induction on $n$. If $n=1$, then $R$ is a finite field, so the statement follows (see for example [lidl, Theorem 6.27]). So, assume $n>1$.
There exists an integer $k$ such that $J(R)^{k+1}=0$ and $J(R)^{k}\neq 0$. Define $S=R/J(R)^{k}$. By the induction hypothesis there exist $\overline{\alpha},\overline{\beta}\in S$ such that $\overline{1}+\overline{\alpha}^{2}+\overline{\beta}^{2}=0$ in $S$. Therefore there exist $\alpha,\beta\in R$ and $w\in J(R)^{k}$ such that $1+\alpha^{2}+\beta^{2}=w$. This implies that at least one of $\alpha,\beta$ is not in $J(R)$. We can assume without any loss of generality that $\alpha\notin J(R)$, so $\alpha\in U(R)$. Since $2\in U(R)$, we can define $\alpha^{\prime}=\alpha-w(2\alpha)^{-1}$ and observe that $\alpha^{\prime 2}=\alpha^{2}-w=-1-\beta^{2}$, so $1+\alpha^{\prime 2}+\beta^{2}=0$. ∎
@@ 3. Products of idempotents
In this section we turn to the product of idempotents in a quaternion ring over a finite commutative local principal ring. The restriction to finite commutative local principal rings is dictated by the methods used here. Locality forces idempotents in $M_{2}(R)$ to be conjugate to standard diagonal ones, which makes products of idempotents amenable to orbit considerations. The principal ideal assumption gives a uniform description $J(R)=(x^{k})$, allowing us to parametrize conjugacy classes $\mathcal{O}_{M(a,b)}$ effectively and to evaluate their sizes explicitly, which is essential for the counting formula in Theorem 3.8. Without principality (or without locality), the structure of the powers of the Jacobson radical and the resulting orbit stratification can be substantially more complicated.
We begin by determining the number of idempotent elements.
@@
Suppose firstly that $p>2$. By Theorem 2.4, we know that $H(R)\simeq M_{2}(R)$. Lemmas 2.2 and 2.3 from [dolzan] now give us $|I(H(R))|=2+q^{3n-2}(q^{2}-1)$. On the other hand, if $p=2$, 22 is a zero-divisor in $R$, so $H(R)$ is a finite local ring by [cherdol, Lemma 3.1] and as such contains no nontrivial idempotents. ∎
@@
We may therefore assume that $2\in R$ is invertible. Then we have $H(R)\simeq M_{2}(R)$ by Theorem 2.4, so we have to prove that any $A\in M_{2}(R)$ that is a product of $r$ idempotent matrices for some $r\geq 3$, can also be written as a product of two idempotents. Therefore, assume that $A=E_{1}E_{2}\ldots E_{r}$ for idempotent matrices $E_{1},E_{2},\ldots,E_{r}$. We can also assume without any loss of generality that $E_{1}$ is a nontrivial idempotent. By Lemma 2.1, there exists an invertible matrix $P\in M_{2}(R)$ such that $PE_{1}P^{-1}=M(1,0)$. Thus $PAP^{-1}=M(1,0)F_{2}\ldots F_{r}$ for some (idempotent) matrices $F_{2},\ldots,F_{r}$. This implies that $PAP^{-1}=M(a,b)$ for some $a,b\in R$. By Lemma 3.3, this implies that $A$ is a product of two idempotents. ∎
@@
$\left|\mathcal{O}_{M(a,b)}\right|=\begin{cases}q^{2n-k-1}(q^{2}-1);\text{ if }k<l\leq n,\\ q^{2n-2l-1}(q+1);\text{ if }a\neq 0,b=0,\\ 1;\text{ if }a=b=0.\end{cases}$
@@
We know that $\left|\mathcal{O}_{M(a,b)}\right|=\frac{|GL_{2}(R)|}{|\mathrm{Stab}(M(a,b))|}$ and that $|GL_{2}(R)|=q^{4n-3}(q-1)(q^{2}-1)$ (see for example [dolzan, Lemma 2.3]), so we have to calculate the sizes of the respective stabilizers.
Let us calculate $|\mathrm{Stab}(M(a,0))|$ for $a\neq 0$. Observe that $P=\left(\begin{array}[]{cc}\alpha&\beta\\ \gamma&\delta\end{array}\right)\in\mathrm{Stab}(M(a,0))$ if and only if $\gamma a=\beta a=0$, so $\beta,\gamma\in J(R)^{n-l}$. Therefore, we have $q^{2l}$ possibilities for choosing $\beta$ and $\gamma$. Since $a\neq 0$, we have $\beta,\gamma\in J(R)$, so $\alpha,\delta\in U(R)$. This gives us $|\mathrm{Stab}(M(a,0))|=q^{2(l+n-1)}(q-1)^{2}$ and therefore $\left|\mathcal{O}_{M(a,0)}\right|=q^{2n-2l-1}(q+1)$.
Let us finally calculate $|\mathrm{Stab}(M(a,b))|$ in the case $k<l\leq n$. Observe that $P=\left(\begin{array}[]{cc}\alpha&\beta\\ \gamma&\delta\end{array}\right)\in\mathrm{Stab}(M(a,b))$ if and only if $PM(a,b)=M(a,b)P$ if and only if $\gamma a=\gamma b=0$ and $a\beta=(\alpha-\delta)b$. Since $k<l$, the conditions $\gamma a=\gamma b=0$ are equivalent to the fact that $\gamma\in J(R)^{n-k}$. Since $|J(R)^{n-k}|=q^{k}$, we have $q^{k}$ possibilities for choosing $\gamma$. By Lemma 2.2, there also exists $x\in J(R)$ such that $J(R)^{t}=(x^{t})$ for every $t=0,1,\ldots,n$. Therefore $a=ux^{l}$ and $b=vx^{k}$ for some $u,v\in U(R)$. Since $\gamma\in J(R)$, the fact that $P$ is an invertible matrix implies that $\alpha,\delta\in U(R)$. Choose any $\beta\in R$ and any $\alpha\in U(R)$, giving us $q^{n}$ and $q^{n-1}(q-1)$ possibilities, respectively. The condition $a\beta=(\alpha-\delta)b$ can be rewritten as $v^{-1}u\beta x^{l}=(\alpha-\delta)x^{k}$. Denote $v^{-1}u\beta=\beta_{0}+\beta_{1}x+\ldots+\beta_{n-1}x^{n-1}$ and $\alpha-\delta=\epsilon_{0}+\epsilon_{1}x+\ldots+\epsilon_{n-1}x^{n-1}$. This implies $\epsilon_{0}=\epsilon_{1}=\ldots=\epsilon_{l-k-1}=0$ and $\epsilon_{l-k}=\beta_{0},\epsilon_{l-k+1}=\beta_{1},\ldots,\epsilon_{n-k-1}=\beta_{n-l}$. (Note here that $k<l$ implies that we always have $\epsilon_{0}=0$, so $\delta$ is automatically invertible if $\alpha$ is invertible.) Therefore, the $k$ elements $\epsilon_{n-k},\epsilon_{n-k+1},\ldots,\epsilon_{n-1}$ can be chosen arbitrarily (giving us $q$ choices for each one, since $J(R)^{m}/J(R)^{m+1}$ is a one-dimensional vector space over $R/J(R)$ by Lemma 2.2). Therefore, we have $|\mathrm{Stab}(M(a,b))|=q^{k}q^{n}q^{n-1}(q-1)q^{k}=q^{2k+2n-1}(q-1)$ and thus $\left|\mathcal{O}_{M(a,b)}\right|=q^{2(n-k-1)}(q^{2}-1)$. ∎
@@
We can therefore assume that $2\in R$ is invertible. Then we have $H(R)\simeq M_{2}(R)$ by Theorem 2.4, so Lemma 3.3 and Theorem 3.4 tell us that $|Z|=|\bigcup_{a,b\in R}\mathcal{O}_{M(a,b)}|+1$. Again using Lemma 2.2, we know there exists $x\in J(R)$ such that $J(R)^{t}=(x^{t})$ for every $t=0,1,\ldots,n$. Suppose $a\in J(R)^{l}\setminus J(R)^{l+1}$ and $b\in J(R)^{k}\setminus J(R)^{k+1}$ for some integer $0\leq l\leq n-1$ and some integer $0\leq k\leq n$. Using Corollary 3.7, we therefore have $\mathcal{O}_{M(a,b)}=\mathcal{O}_{M(a,0)}$ if $k\geq l$ and $\mathcal{O}_{M(a,b)}=\mathcal{O}_{M(a,x^{k})}$ if $k<l$. This gives us
@@
$|Z|=2+\sum_{k=0}^{n-1}q^{2(n-k-1)}(q^{2}-1)+\\ \sum_{l=0}^{n-1}\left(q^{n-1-l}(q-1)\left(q^{2n-2l-1}(q+1)+\sum_{k=0}^{l-1}q^{2(n-k-1)}(q^{2}-1)\right)\right).$
@@
$|Z|=2+(q^{2n}-1)+\frac{q(q^{3n}-1)(q^{2}-1)}{q^{3}-1}+\sum_{l=0}^{n-1}(q-1)q^{3(n-l)-1}(q^{2l}-1)=\\ 1+q^{2n}+\frac{q(q^{3n}-1)(q^{2}-1)}{q^{3}-1}+q^{3n}-q^{2n}-\frac{q^{2}(q-1)(q^{3n}-1)}{q^{3}-1},$
@@
$\frac{10+16\alpha^{3}}{13}$